# Quantum Critical Collapse Abhors a Naked Singularity

### v1

@@ Abstract
Since all non-spherical modes decay except a single spherical growing mode, it is natural to focus exclusively on the $s$-wave sector, which captures the dominant quantum corrections. For a minimally coupled, massless scalar field $f$ in $D=d+1$ dimensions, one isolates the $s$-wave via spherical reduction to two dimensions with a dilaton $\phi$ [30]:
@@
Here $R$ reproduces the conformal anomaly of a free scalar, while the dilaton terms encode the effects of spherical reduction, ensuring that the essential quantum imprint of the higher-dimensional theory is captured. Crucially, the universal dilaton coupling makes the anomaly structure extremely general across different matter models.
@@ Acknowledgements.
We thank David Garfinkle, Viqar Husain, Gabor Kunstatter, and Edward Witten for useful discussions. In particular, we thank Roberto Emparan, Carsten Gundlach, and Gustavo J. Turiaci for reading the draft and providing valuable feedback that clarified various subtleties. CHW also thanks Gustavo J. Turiaci for numerous discussions that significantly improved this work. He further acknowledges an audience member at the 39th Pacific Coast Gravity Meeting, whose suggestion inspired this project, and is deeply indebted to Gary Horowitz and Jiuci Xu for their early collaboration and many insightful discussions. CHW is supported by the University of Washington and DOE Award DE-SC0011637. MT is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 852386). MT is also supported by the Emmy Noether Fellowship program at the Perimeter Institute for Theoretical Physics.

### v2

@@ Abstract
Since all non-spherical modes decay except for a single spherical growing mode, it is sufficient to focus on the $s$-wave sector in describing the infrared dynamics. For a minimally coupled, massless scalar field $f$ in $D=d+1$ dimensions, one isolates the $s$-wave via spherical reduction to two dimensions with a dilaton $\phi$ [30]:
@@
Here $R$ reproduces the conformal anomaly of a free scalar. The additional dilaton-anomaly terms preserve the state-independent ultraviolet structure of the higher-dimensional theory, while the infrared dynamics are governed solely by the $s$-wave, since no quantum state can excite the decaying higher angular modes. Consequently, the formulation captures all dynamically relevant quantum effects and retains the essential quantum imprint of the full higher-dimensional theory. Crucially, the universal dilaton coupling renders the anomaly structure broadly applicable across different matter models.
@@ Acknowledgements.
We thank David Garfinkle, Viqar Husain, Gabor Kunstatter, and Edward Witten for useful discussions. In particular, we thank Roberto Emparan, Carsten Gundlach, and Gustavo J. Turiaci for reading the draft and providing valuable feedback that clarified various subtleties. CHW also thanks Gustavo J. Turiaci for numerous discussions that significantly improved this work. He further acknowledges an audience member at the 39th Pacific Coast Gravity Meeting, whose suggestion inspired this project, and is deeply indebted to Gary Horowitz and Jiuci Xu for their early collaboration and many insightful discussions. CHW is supported by the University of Washington and DOE Award DE-SC0011637 and DE-SC0026287. MT is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 852386). MT is also supported by the Emmy Noether Fellowship program at the Perimeter Institute for Theoretical Physics.